OpenStudy (anonymous):

What is the value of the expression 10 − (fraction 1 over 2)4 ⋅ 48?

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

\[10-(\frac{1}{2})^4 \cdot 48\] is this your equation?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

ok so order of operations is coming into play PEMDAS P is the parenthesis and E is the exponent... looks we got a 2 for 1 special solve \[(\frac{1}{2})^4 \] first the exponent is 4 which means write \[\frac{1}{2}\] 4 times

OpenStudy (anonymous):

that is 2 wholes

OpenStudy (usukidoll):

\[\frac{1}{2}\frac{1}{2}\frac{1}{2}\frac{1}{2}\] no we are doing multiplication

OpenStudy (anonymous):

k

OpenStudy (usukidoll):

\[\frac{1x1x1x1}{2x2x2x2} \] what's 1 x 1 x 1 x 1 what's 2 x 2 x 2 x 2

OpenStudy (anonymous):

1 x 1 x 1 x 1=1 2 x 2 x 2 x 2=16

OpenStudy (usukidoll):

yup so our fraction is 1/16 \[10-\frac{1}{16} \cdot 48\] now we have M D which is multiplication and division what is \[\frac{1}{16} \cdot 48\]

OpenStudy (anonymous):

3?

OpenStudy (usukidoll):

very good :) now we just have 10-3 we don't have A (addition) so skip it We do have S (subtraction) 10 - 3 = ?

OpenStudy (anonymous):

7

OpenStudy (usukidoll):

yes. correct!

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

r u still up for more questions

OpenStudy (usukidoll):

I'm a bit busy with something. taking a survey.

OpenStudy (anonymous):

ok so no more??? :(

OpenStudy (usukidoll):

I'll come back.. it will only take 10 more minutes.

OpenStudy (anonymous):

ok